In 2005, Dan raced his car 500 miles. His rate was 157.603 mph. What was his time (to the nearest thousandth of an hour)?
Answer by tutorcecilia(2152)

Use the distance formula: Distance = (Rate)(Time)
500 miles = (157.603)(t) [plug-in the values]
500/157.603=157.603t/157.603 [isolate the t-term]
3.172528=t
or
t=3.173 hours
.
Check by plugging all of the variables back into the original equation and solve:
D=rt
500=(157.603)(3.173)
500=500.074 [close enough]